OpenStudy (anonymous):

solve the system of equation by graphing x+y=5 x-y=-3

hero (hero):

y = 5 - x y = -3 + x y = y 5-x = 3 +x 2 = 2x x = 1

OpenStudy (amistre64):

y = 5-x y = -3-x

hero (hero):

Amistre, I made your mistake the first time as well

OpenStudy (amistre64):

:) im contagious lol

OpenStudy (anonymous):

dont know how to do this problem and how to graph it as well, im taking that the answer is y=5-x y=-3-x am i right>?

OpenStudy (amistre64):

doh!! ....

OpenStudy (anonymous):

how do you graph it

OpenStudy (amistre64):

-y = -3-x y = 3+x .... murmur

myininaya (myininaya):

we have x+y=5 and x-y=-3 add now (x+x)+(y-y)=(5-3) 2x+0=2 x=1 => 1+y=5 =>y=4

hero (hero):

The two equations where it says y = you graph those

OpenStudy (amistre64):

graphing is rather self explanatory; you create a table of points and connect the dots

OpenStudy (anonymous):